# Why Smishing Works

### v2

@@ 5.2.5. Occupation
Likely to interact: The post hoc test results show that the average likely interaction rates with fake messages was significantly lower for students when compared to art, writing, or journalism (p = 0.048, 95% C.I. = [-2.38, -0.01]), business, management, or financial (p $<$ 0.001, 95% C.I. = [-2.38, -0.45]), IT (p = 0.004, 95% C.I. = [-2.59, -0.30]), and service occupations (p = 0.012, 95% C.I. = [-2.77, -0.21]).
Comfortability with interacting: The post hoc test results show that the average comfort levels with fake messages was significantly lower for students when compared to IT (p = 0.010, 95% C.I. = [-1.96, -0.16]), and service occupations (p = 0.022, 95% C.I. = [-2.00,-0.09]).
Accuracy: The post hoc test results for average accuracy when identifying fake messages were not significant.
@@
Real Messages Sender 313 -.190 .192 .827 [.621 - 1.100]
Entity 316 .369 .013 1.446 [1.080 - 1.935]
URL 1096 -.078 .485 .925 [.742 - 1.152]
CTA 392 .549 ¡.001 1.732 [1.330 - 2.255]
Scenario 531 .074 .538 1.077 [.850 - 1.365]
Const. -.334 ¡.001 .716 -
Fake Messages Sender 238 .463 .012 1.590 [1.108 - 2.280]
Entity 269 -.117 .481 .889 [.642 - 1.233]
URL 660 .490 ¡.001 1.632 [1.242 - 2.145]
CTA 166 -.227 .103 .797 [.607 - 1.047]
Scenario 505 -.007 .970 .993 [.671 - 1.468]
Const. .417 ¡.001 1.517 -
All Messages Sender 551 .115 .281 1.122 [.910 - 1.385]
Entity 585 .209 .053 1.232 [.997 - 1.522]
URL 1756 .102 .229 1.107 [.938 - 1.307]
CTA 558 .294 .007 1.342 [1.083 - 1.661]
Scenario 1036 .026 .767 1.026 [.865 - 1.218]
Const. -.052 .477 .949 -
@@ 5.3.1. Areas of Interest and their relationship with accurately identifying real and fake messages
To explore the impact of message attributes on a user’s likelihood to interact, comfort levels, and accuracy in identifying real and fake messages, we implemented a task where participants clicked on pictures of SMS messages to indicate which part of the image influenced their feelings about interacting. Participant clicks were then categorized into five distinct AOI: the sender, entity mentioned in the message, the URL contained in the message, the call to action in the message, and the text describing the scenario of the message. An exception was made for one of our fake messages that lacked a URL, entity, call to action, or scenario, leading to its exclusion from the AOI analysis. In total, participants generated 10,156 clicks across 214 participants and 16 images, with 5,466 of these clicks landing within one of the specified AOIs. 980 clicks resulted in an ”i don’t know” decision, which we separate from this analysis. The remaining 4486 clicks lead to a ’real’ or ’fake’ decision. The messages with their corresponding AOI and user clicks can be found in Figure LABEL:fig:aoi_images in the appendix.
@@
Real Messages Clicking on the entity increased the odds of correctly identifying real messages by 44.6% (95% C.I. [1.08, 1.935]), and clicking on the call to action increased the odds by 73.2% (95% C.I. [1.33, 2.26]).
Fake Messages Clicking on the sender increased the odds of correctly identifying fake messages by 58.9% (95% C.I. [1.08, 2.28]), while clicking on the URL increased the odds by 63.2% (95% C.I. [1.24, 2.15]).
All Messages Clicking on the call to action increased the odds of correctly identifying messages by 34.2% (95% C.I. [1.08, 1.66]).
@@
Real Meassages Sender -.24 .001 [-.39, -.09] -.204 .036 [-.40, -.01]
Entity .028 .708 [-.12, .17] .039 .682 [-.15, .23]
URL -.032 .576 [-.144, .080] .037 .616 [-.11, .18]
CTA .216 .002 [.082, .35] .194 .029 [.02, .37]
Scenario .106 .084 [-.014, .23] .089 .260 [-.07, .25]
Constant 2.69 <.001 [2.60, 2.79] 2.84 <.001 [2.71, 2.96]
Fake Messages Sender -.44 <.001 [-.61, -.26] -.562 <.001 [-.79, -.33]
Entity .003 .973 [-.16, .17] .004 .974 [-.21, .22]
URL -.273 <.001 [-.41, -.14] -.302 <.001 [-.48, -.13]
CTA -.054 .591 [-.25, .14] .053 .687 [-.20, .31]
Scenario .122 .082 [-.02, .26] .207 .024 [.03, .39]
Constant 2.69 <.001 [2.56, 2.81] 2.96 <.001 [2.81, 3.12]
Table 3. Linear Regression results for the effects of AOI on Comfort Levels and Likely Interaction Rates in real and fake messages.
@@
Interaction Rates With Real Messages Clicking on the sender decreased the average likelihood of interacting with a real message by 0.2 points(95% C.I. [-.40,-.01]). On the other hand, clicking on the call to action lead to an increase in the average likelihood of interacting with a real message by 0.19 points(95% C.I. [.02,.37]).
Interaction Rates With Fake Messages Clicking on the sender decreased the likelihood of interacting with a fake message by -0.6 points(95% C.I. [-.79,-.33]). Clicking on the URL also decreased the average likelihood of interacting with a fake message by -0.3 points(95% C.I. [-.48,-.13]). Alternatively, clicking on the scenario lead to an increase in the average likelihood of interacting with a fake message by 0.21 points(95% C.I. [.03,.39]).
Comfort Levels With Real Messages Clicking on the sender decreased the average comfort level interacting with a real message by 0.24 points(95% C.I. [-.39,-.09]). On the other hand, clicking on the call to action lead to an increase in the likelihood of interacting with a real message by 0.22 points(95% C.I. [.08, .35]).
Comfort Levels With Fake Messages Clicking on the sender decreased the likelihood of interacting with a fake message by -0.44 points(95% C.I. [-.61,-.26]). Clicking on the URL also decreased the likelihood of interacting with a fake message by -0.27 points(95% C.I. [-.41,-.14]).
@@
Very Comf (C/I) 20.41/1.97 21.24/3.67 19.94/2.23 22.36/4.76 21.31/4.25 0.64/12.24 0.00/18.67 2.18/12.33 1.16/13.04 3.33/17.76
Somewhat Comf (C/I) 32.65/7.24 35.40/6.42 36.42/8.48 37.27/11.90 32.79/12.26 6.41/36.73 12.50/36.00 7.90/34.93 9.30/36.96 9.58/40.79
Neither (C/I) 18.37/12.50 19.47/21.10 26.01/20.54 20.50/19.05 22.95/21.23 12.82/22.45 13.97/20.00 15.26/29.45 12.79/34.78 11.25/23.68
Somewhat Uncomf (C/I) 21.43/31.58 14.16/24.77 12.72/28.13 15.53/26.19 16.94/28.30 19.87/18.37 19.85/14.67 22.07/13.70 24.42/15.22 27.92/13.16
Very Uncomf (C/I) 7.14/46.71 9.73/44.04 4.91/40.63 4.35/38.10 6.01/33.96 60.26/10.20 53.68/10.67 52.59/9.59 52.33/0.00 47.92/4.61
@@
Definitely (C/I) 20.20/2.63 17.70/5.50 17.97/3.13 22.50/3.94 18.13/3.27 1.27/21.57 0.73/25.00 1.36/17.81 1.16/15.22 2.06/22.73
Very Probably (C/I) 21.21/3.29 26.55/3.67 26.38/4.91 19.38/3.15 28.02/5.61 3.18/25.49 6.57/15.79 6.54/23.29 9.30/26.09 7.41/24.68
Probably (C/I) 20.20/5.92 15.04/6.42 18.26/8.04 23.13/11.02 15.38/7.94 2.55/11.76 7.30/19.74 7.90/18.49 6.98/19.57 7.41/19.48
Possibly (C/I) 17.17/13.16 22.12/11.93 19.13/15.18 18.13/17.32 20.88/17.29 6.37/33.33 16.89/25.00 10.00/23.29 17.44/30.43 12.34/19.48
Probably Not (C/I) 15.15/23.68 10.62/21.10 11.30/23.66 10.63/22.05 10.44/21.96 19.11/3.92 20.44/3.95 17.44/4.11 24.42/10.87 24.69/5.19
Definitely Not (C/I) 6.06/51.32 7.96/51.38 6.96/45.09 6.25/42.52 7.14/43.93 67.52/3.92 55.47/7.89 57.49/7.53 51.16/2.17 48.15/5.84
@@ 5.4. Answering RQ:4
We examined the correlation between post-survey results regarding user behavior, interaction likelihood, comfort levels, and accuracy in identifying real and fake messages using Spearman’s rank correlation coefficient. These post-survey questions comprised Likert scales focused on privacy and security, totaling five in number.
@@
First, we examine the relationship between SA-6 scores and accuracy and interaction rates. According to our Spearman’s correlation analysis, we did not observe any significance between accuracy or interaction rates and SA-6 scores. Moving on, we investigate the relationship between IUIPC scores and our accuracy and interaction metrics. IUIPC scores measure the user’s internet privacy concerns. After analyzing Spearman’s correlation, we found a moderate, statistically significant decrease in interaction rates with fake messages with an increase in IUIPC scores (${}_{cor}$ = -0.303, p $>$ 0.001). We also found a small, statistically significant decrease in interact rates with real messages with an increase in IUIPC scores (${}_{cor}$ = -0.206, p $=$ 0.002). These findings suggest that participants who expressed higher concerns about their internet privacy were less likely to interact with both real and fake messages. However, no significant correlation was found between IUIPC scores and the overall message accuracy rates.
@@
We then explore the relationship between RSeBIS scores and our accuracy and interaction metrics. RSeBIS is a refined version of the Security Behavior Intention Scale (SeBIS) that measures the end user’s security behaviors. Our Spearman’s correlation revealed a small, statistically significant decrease in the interaction rates with fake messages as RSeBIS scores increased (${}_{cor}$ = -0.237, p ¡ 0.001). These results indicate that participants who scored higher on the RSeBIS scale reported being less likely to interact with our fake messages. However, we found no significant correlations between RSeBIS scores and overall accuracy or interaction rates with real messages.
@@
In our next post-survey test, we examined the correlation between familiarity with Security Terms and overall accuracy and interaction rates. Our Security Term Familiarity test measured participants’ familiarity with security terms related to SMS messages. Our analysis of the Spearman’s correlation found a small, statistically significant decrease in interaction rates with fake messages with higher average Security Term Familiarity (${}_{cor}$ = -0.210, p = 0.002). However, we found no significant correlations between Security Term Familiarity and overall accuracy or interaction rates with real messages.
@@
Finally, we investigated the relationship between the Attention Control (ATTC) Scale and participants’ accuracy and interaction rates. The ATTC scale measures individual differences in attention control based on the factors of attention focus and attention shifting. Our Spearman’s correlation results show a small, statistically significant decrease in the interaction rates with both fake messages (${}_{cor}$ = -0.235, p = 0.001) and real messages (${}_{cor}$ = -0.231, p = 0.001) in accordance with a higher ATTC scale score. These results indicate that participants with a higher ATTC scale score were less likely to interact with either fake or real messages. However, we found no significant correlations between ATTC scale scores and overall accuracy.
@@ 6. Discussion
In their work, Rahman et al. (rahman2023users) conducted a user study on smishing to investigate the rate at which users fall for smishing attacks. Their study reported that 16.92% of participants fell for a smishing attack. In contrast, our study found that 34.4% of fake messages were incorrectly identified as real. The difference in our rates can be attributed to several factors. In their study, successful smishing attacks were counted only when users actively interacted with the phishing messages. Thus, users who believed the messages were real but did not interact with them were not counted as being tricked by the attack. In our experiment, we created a controlled environment where users were asked to decide whether they believed the message was real or fake. Additionally, in this controlled environment, users did not benefit from potential built-in SMS protections or anti-phishing applications, such as security indicators in SMS apps that notify users when a message comes from an unknown number. Given these differences, we believe that our results offer a unique perspective on how users may be tricked by smishing.
@@ 6.4. Recommendations
Mobile Users need more education on short codes. Our analysis revealed that while participants clicked 313 times on the sender in real messages, this led to (152/313) of those messages being incorrectly identified as fake and (62/313) caused uncertainty. Notably, messages with short codes showed lower average comfort levels and likely interaction rates among participants. While short codes have been shown to be able to be spoofed, they are on average safe than regular phone numbers.
Brands should do more to make their links recognizable. We recommend that brands use clear and easy to associate links with their business. In our testing, a real Microsoft password reset SMS with a unique link shortener (aka.ms) had an average accuracy of 34.3%. Over half of participants (114/214) identified aka.ms, as causing them to feel less comfortable with the SMS. For users unfamiliar with Microsoft’s ownership of aka.ms, such shortened links may seem ambiguous and suspicious. Among those who identified the URL, 69.2% incorrectly marked it as fake.
How user training and app developers could improve user detection.
Our findings can contribute to the design of improved educational tools and warning systems for detecting smishing attempts. By recognizing specific AOI that are often misunderstood as fraudulent or overlooked by users, developers of anti-smishing technologies can target these elements directly. This is supported by the findings that users with higher attention levels were more adept at correctly identifying smishing messages. In educational training, simulated scenarios akin to those utilized in this study could be crafted to underscore the significance of scrutinizing sender information and URLs in messages, emphasizing their pivotal role as primary indicators of a fake message. Furthermore, similar to how website developers can draw attention to particular aspects of web pages, mobile messaging app developers can enhance users’ ability to identify fake messages by directing their attention to these areas, which may make users less comfortable or inclined to interact with fake messages.

### v4

@@ 5.2.6. Occupation
Likely to interact: The post hoc test results show that the average likely interaction rates with fake messages was significantly lower for students when compared to business, management, or financial (p 0 0.004, 95% C.I. = [-2.32, -0.29]) and IT (p = 0.008, 95% C.I. = [-2.70, -0.25]).
Comfortability with interacting: The post hoc test results show that the average comfort levels with fake messages was significantly lower for students when compared to IT (p = 0.009, 95% C.I. = [-1.98, -0.17]), and service occupations (p = 0.031, 95% C.I. = [-2.15,-0.06]).
Accuracy; We found no statistically significant difference when considering the criteria of accuracy.
RQ2 Summary: We found that differences in the Sex, Ethnicity, Age, Income and Occupation of participants had a significant impact on the user’s likely interaction rates, comfort levels and accuracy in identifying the messages.
@@
Real Messages Sender 304 -.014 .925 .986 [.74 - 1.31]
Entity 294 .216 .117 1.241 [.95 - 1.62]
URL 1316 -.050 .646 .951 [.77 - 1.18]
CTA 378 .395 .002 1.484 [1.16 - 1.90]
Scenario 573 .101 .379 1.106 [.88 - 1.39]
Const. - -.875 <.001 .417 -
Fake Messages Sender 239 .671 <.001 1.956 [1.42 - 2.69]
Entity 278 -.043 .770 .958 [.72 - 1.28]
URL 772 .408 <.001 1.504 [1.18 - 1.92]
CTA 165 .104 .567 1.110 [.78 - 1.58]
Scenario 579 -.103 .412 .902 [.70 - 1.15]
Const. - -.249 .025 .779 -
All Messages Sender 543 .334 .001 1.396 [1.14 - 1.70]
Entity 572 .143 .145 1.154 [.95 - 1.40]
URL 2088 .102 .200 1.107 [.95 - 1.29]
CTA 543 .218 .031 1.244 [1.02 - 1.52]
Scenario 1152 .068 .405 1.071 [.91 - 1.26]
Const. - -.622 <.001 .537 -
@@ 5.3.1. Areas of Interest and their relationship with accurately identifying real and fake messages
To explore the impact of message attributes on a user’s likelihood to interact, comfort levels, and accuracy in identifying real and fake messages, we implemented a task where participants clicked on pictures of SMS messages to indicate which part of the image influenced their feelings about interacting. Participant clicks were then categorized into five distinct AOI: the sender, entity mentioned in the message, the URL contained in the message, the call to action in the message, and the text describing the scenario of the message. An exception was made for one of our fake messages that lacked a URL, entity, call to action, or scenario, leading to its exclusion from the AOI analysis. In total, participants generated 8,832 clicks across 187 participants and 16 images, with 4,898 of these clicks landing within one of the specified AOIs. 706 clicks resulted in an ”i don’t know” decision, which we separate from this analysis. The remaining 4192 clicks lead to a ’real’ or ’fake’ decision. In 81% of the individual message tasks, participant clicked within at least one AOI. The messages with their corresponding AOI and user clicks can be found in Figure 4 in the appendix.
@@
Real Messages Clicking on the CTA increased the odds of correctly identifying real messages by 48.4% (95% C.I. [1.16, 1.90])).
Fake Messages Clicking on the sender increased the odds of correctly identifying fake messages by 95.6% (95% C.I. [1.42, 2.69]), while clicking on the URL increased the odds by 50.4% (95% C.I. [1.18, 1.92]).
All Messages Clicking on the sender increased the odds of correctly identifying all messages by 39.6% (95% C.I. [1.14, 1.70]) while clicking on the CTA increased the odds of correctly identifying messages by 24.4% (95% C.I. [1.02, 1.52]).
@@
Real Meassages Sender -.29 <.001 [-.45,-.14] -0.28 .01 [-.48,-.07]
Entity .07 .404 [-.09,.22] 0.08 .43 [-.12,.28]
URL -.04 .531 [-.16,.08] 0.02 .79 [-.13,.18]
CTA .15 .050 [.00,.29] 0.18 .06 [-.01,.37]
Scenario .12 .065 [-.01,.25] .10 .25 [-.07,.26]
Constant 2.67 <.001 [2.56,2.77] 2.78 <.001 [2.64,2.91]
Fake Messages Sender -.45 <.001 [-.63,-.26] -.56 <.001 [-.80,-.32]
Entity -.03 .763 [-.20,.14] .00 .969 [-.22,.23]
URL -.33 <.001 [-.47,-.18] -.34 <.001 [-.53,-.14]
CTA -.09 .395 [-.30,.12] .03 .841 [-.25,.31]
Scenario .05 .478 [-.09,.20] .17 .085 [-.02,.37]
Constant 2.71 <.001 [2.58,2.85] 2.94 <.001 [2.77,3.12]
Table 3. Linear Regression results of AOI effects on Comfort Levels and Likely Interaction Rates in real and fake messages.
@@
Interaction Rates With Real Messages Clicking on the sender decreased the average likelihood of interacting with a real message by 0.28 points(95% C.I. [-.48,-.07]).
Interaction Rates With Fake Messages Clicking on the sender decreased the likelihood of interacting with a fake message by 0.56 points(95% C.I. [-.79,-.33]). Clicking on the URL also decreased the average likelihood of interacting with a fake message by 0.34 points(95% C.I. [-.53,-.14]).
Comfort Levels With Real Messages Clicking on the sender decreased the average comfort level interacting with a real message by 0.29 points(95% C.I. [-.45,-.14]).
Comfort Levels With Fake Messages Clicking on the sender decreased the likelihood of interacting with a fake message by 0.45 points(95% C.I. [-.63,-.26]). Clicking on the URL also decreased the likelihood of interacting with a fake message by 0.33 points(95% C.I. [-.47,-.18]).
@@
Very Comf (C/I) 16.5/2.1 21.6/3.3 19.2/2.3 20.0/4.9 20.6/3.8 0.0/8.5 0.0/12.8 1.8/7.6 1.3/7.8 2.3/12.2
Somewhat Comf (C/I) 35.3/9.4 35.3/12.5 35.5/10.0 35.6/11.8 33.3/12.3 6.2/28.2 10.9/21.4 7.3/19.6 6.4/29.7 8.9/29.4
Neither (C/I) 17.6/18.3 18.6/28.8 26.7/30.0 22.2/30.5 21.8/31.9 11.6/18.3 14.7/35 14.3/41.2 12.8/32.8 10.7/32.1
Somewhat Uncomf (C/I) 22.4/28.3 14.7/22.3 13.4/25.5 17.0/21.7 18.2/25.2 19.9/31 18.6/19.7 21.6/20.4 24.4/23.4 27.6/19.0
Very Uncomf (C/I) 8.2/41.9 9.8/33.2 5.2/32.3 5.2/31.0 6.1/26.8 62.3/14.1 55.8/11.1 55.0/11.2 55.1/6.3 50.5/7.2
@@
Definitely (C/I) 16.3/1.1 18.6/2.2 18/2.6 24.6/2 18.9/3.5 0.7/15.3 0.8/18.5 0.6/12.4 0.0/15.4 0.9/18.2
Very Probably (C/I) 23.3/4.2 28.4/6.5 27.1/6.1 17.2/6.4 27.4/5.0 3.4/15.3 5.4/11.8 5.6/17.2 7.7/23.1 6.5/18.7
Probably (C/I) 19.8/8.9 13.7/9.8 15.7/10.5 20.9/10.3 15.2/11 2.7/13.9 7.8/21.0 7.9/16 7.7/13.8 7.5/19.1
Possibly (C/I) 17.4/12.6 21.6/17.9 20.3/17.6 17.9/20.6 20.1/20.8 4.1/27.8 7.8/22.7 7.9/26.8 6.4/23.1 9.8/19.6
Probably Not (C/I) 16.3/26.3 9.8/23.4 11.8/24.4 11.9/24 10.4/20.5 19.0/19.4 20.9/13.4 17.3/14.8 23.1/16.9 22.9/15.1
Definitely Not (C/I) 7.0/46.8 7.8/40.2 7.2/38.9 7.5/36.8 7.9/39.1 70.1/8.3 57.4/12.6 60.7/12.8 55.1/7.7 52.3/9.3
@@
RQ3 Summary: Our findings indicate that different Areas of Interest (AOIs) affected the participants’ comfort levels and interaction rates with both real and fake messages differently. Specifically, for fake messages, identifying the sender and URL significantly reduced interaction rates and comfort levels. Conversely, for real messages, engaging with the call to action led to higher interaction rates and increased comfort levels. Furthermore, participants who incorrectly identified an AOI as influencing their comfort were significantly more likely to interact with fake messages. In contrast, those who accurately identified the sender were significantly less likely to engage with real messages.
@@ 5.4. Answering RQ:4
We examined the correlation between post-survey results and user behavior, interaction rates, comfort levels, and accuracy in identifying real and fake messages using Spearman’s rank correlation coefficient. These post-survey questions comprised Likert scales focused on privacy and security.
First, we examine the relationship between SA-6 scores and our accuracy and interaction rates. According to our Spearman’s correlation analysis, we did not observe any significance between accuracy or interaction rates and SA-6 scores.
@@
Moving on, we investigate the relationship between IUIPC scores and our accuracy and interaction rates. After analyzing Spearman’s correlation, we found a moderate, statistically significant decrease in interaction rates with fake messages with an increase in IUIPC scores (rcor = -0.337, p $<$ 0.001). We also found a small, statistically significant decrease in interact rates with real messages with an increase in IUIPC scores (rcor = -0.260, p $<$ 0.001). These findings suggest that participants who expressed higher concerns about their internet privacy were less likely to interact with both real and fake messages. However, no significant correlation was found between IUIPC scores and the overall message accuracy rates.
@@
We then explore the relationship between RSeBIS scores and our accuracy and interaction metrics. Our Spearman’s correlation revealed a small, statistically significant decrease in the interaction rates with fake messages as RSeBIS scores increased (rcor = -0.237, p $<$ 0.001). These results indicate that participants who scored higher on the RSeBIS scale reported being less likely to interact with our fake messages. However, we found no significant correlations between RSeBIS scores and overall accuracy or interaction rates with real messages.
@@
In our next post-survey test, we examined the correlation between familiarity with Security Terms and overall accuracy and interaction rates. Our analysis of the Spearman’s correlation found a small, statistically significant decrease in interaction rates with fake messages with higher average Security Term Familiarity (rcor = -0.175, p = 0.017). However, we found no significant correlations between Security Term Familiarity and overall accuracy or interaction rates with real messages.
@@
Finally, we investigated the relationship between the Attention Control (ATTC) Scale and participants’ accuracy and interaction rates. Our Spearman’s correlation results show a small, statistically significant decrease in the interaction rates with both fake messages (rcor = -0.247, p $<$ 0.001) and real messages (rcor = -0.238, p $<$ 0.001) in accordance with a higher ATTC scale score. These results indicate that participants with a higher ATTC scale score were less likely to interact with either fake or real messages. However, we found no significant correlations between ATTC scale scores and overall accuracy.
RQ4 Summary: Our results indicate that participants scoring higher on the IUIPC, RSeBIS, Security Term Familiarity, and ATTC scales were less likely to engage with fraudulent messages. However, none of the behavioral scales significantly enhanced overall message accuracy. Consequently, individuals more concerned about privacy and exhibiting greater attention control appeared to be more skeptical towards all messages, leading them to be less comfortable and less likely to interact.
@@ 6. Discussion
In their work, Rahman et al. (rahman2023users, ) conducted a user study on smishing to investigate the rate at which users fall for smishing attacks. Their study reported that 16.92% of participants fell for a smishing attack. In contrast, our study found that 34.4% of fake messages were incorrectly identified as real. The difference in our rates can be attributed to several factors. In their study, successful smishing attacks were counted only when users actively interacted with the phishing messages. Thus, users who believed the messages were real but did not interact with them were not counted as being tricked by the attack. In our experiment, we created a controlled environment where users were asked to decide whether they believed the message was real or fake. Additionally, in this controlled environment, users did not benefit from potential built-in SMS protections or anti-phishing applications, such as security indicators in SMS apps that notify users when a message comes from an unknown number. Given these differences, we believe that our results offer a unique perspective on how users may be tricked by smishing.
@@ 6.4. Recommendations
Mobile Users need more education on short codes. Our analysis revealed that while participants clicked 313 times on the sender in real messages, this led to (152/313) of those messages being incorrectly identified as fake and another (62/313) caused uncertainty. Notably, messages with short codes showed lower average comfort levels and likely interaction rates among participants. While short codes have been shown to be able to be spoofed, they are more difficult to spoof than regular phone numbers, and are generally considered safer (shortcodespoofing, ).
How user training and app developers could improve user detection. Our findings can contribute to the design of improved educational tools and warning systems for detecting smishing attempts. By recognizing specific AOI that are often focus on or overlooked by users, developers of anti-smishing technologies can improve visibility of elements that lead to better recognition of smishing. This is supported by the findings that users with higher attention levels were more adept at correctly identifying smishing messages. In educational training, simulated scenarios akin to those utilized in this study could be created to underscore the significance of scrutinizing sender information and URLs in messages, emphasizing their pivotal role as primary indicators of a fake message. Furthermore, similar to how website developers can draw attention to particular aspects of web pages, mobile messaging app developers can enhance users’ ability to identify fake messages by drawing their attention to these areas, which may make users less comfortable or inclined to interact with fake messages.